OpenStudy (anonymous):

Help with three problems!! The graph below shows the relationship between distance and time for a car driven at a constant speed. 3. What is the speed? a. 60 mi/h b. 50 mi/h c. 80 mi/h d. 40 mi/h 4. Is this a function relationship? a. yes b. no c. There is not enough information to determine whether the relationship is a function. 5. If this is a function, select a rule to represent it. a. d = t + 50 b. d = 60t c. d = 40t + 10 d. The relationship is not a function.

OpenStudy (anonymous):

I need help

OpenStudy (dumbcow):

d = r*t --> r = d/t to find speed of car divide distance traveled by time

OpenStudy (anonymous):

Still lost

OpenStudy (dumbcow):

what does graph look like?

OpenStudy (anonymous):

Cant upload for some reason

sam (.sam.):

Take the gradient of the graph and you'll get the speed